I am enclosing two pics of my problem ( you do not show how you got the answer there are at least 3 more steps to getting both marked sections.) help2 := Chapter 6, Problem 11P 8 Bookmarks Show all steps: ON Step 4 of 4 A Consider the equilibrium of the block along the vert2 := Chapter 6, Problem 11P 8 Bookmarks Show all steps: ON Step 4 of 4 A Consider the equilibrium of the block along the vert
Answer #1

- 200 - 0.36480 1819 1PX sinis to. 81919x cos 15 = 0 - 200 - 0.07U6P + 0.7911P = 0 -200 + 60.71394P = o. - 200 = -0.71394 P P